40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1335     Entry time: Thu Mar 3 00:59:29 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: upgrade continues 

 I made some more progress on the gyro upgrade today:

  • I found some LM317s at the EE stockroom and put one on the second PD board. I also stuffed it with everything else we need besides the Cougar AP1053. It looks like we won't have one for a while, so I bypassed the Cougar footprint with a wire. I also borrowed another Perkin-Elmer C30642 from Frank/Peter, so this board is now ready for testing at the 40m tomorrow morning.
  • I made some more SMP -> SMA cables for inside the PD boxes:


  • I also borrowed a 9-pin D-sub breakout board from the 40m, which enabled me to use the first PD in its fully assembled form (with the exception of the BNC for DC out---for this we will need to connect the bare leads of a BNC feedthrough to an SMP connector on the board). The gyro cavity is currently locked in the CCW direction with PD S/N 01, and the error, control, TRANS DC and REFL DC are all being sent to the DAQ (the DC out of the PD is being monitored using the breakout board). Some shots:



Signals -- Yellow: REFL DC, Blue: ERR, Magenta: CTRL, Green: TRANS DC

  • I got started with building the temporary breadboard version of the PDH2 servo. It should be done by tomorrow.
  • Tomorrow:
    • Take PD S/N 02 to the 40m and take a transfer function, tune it, and put it in its box
    • Finish building the PDH2 servo, take TF and noise spectrum
    • Install CW MMT and isolation optics
    • Lock cavity in both directions
    • Get new spectrum
ELOG V3.1.3-